OpenStudy (anonymous):

PLEASE HELP!!! the Spanish club is selling churros as a fundraiser. If its goal is to raise at least $150, how many churros must it sell at $1.25 each to meet that goal? write and solve an inequality.

OpenStudy (anonymous):

\[1.25c \ge 150; c \ge 120\]

OpenStudy (anonymous):

\[1.25c \ge 150; c \ge 100\]

OpenStudy (anonymous):

1.25c >= 150 c >= 150/1.25 c >= 120

OpenStudy (anonymous):

\[150c \ge 1.25; c \ge 120\]

OpenStudy (anonymous):

your 1st answer choice that you posted

OpenStudy (anonymous):

The first one? I have two that have c>120....the third and first.

OpenStudy (anonymous):

the first one is correct

OpenStudy (anonymous):

Okay! Thank you soo much!

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

can you help me with another please?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Thank you! Jackie scored 82 and 88 on her first two quizzes. Write and solve a compound inequality to find the possible values for a third quiz score that would give her an average between 80 and 85, inclusive.

OpenStudy (anonymous):

80< 82+88+n< 85; 70 <n <85 3 The 82+88+n is a faction over 3

OpenStudy (anonymous):

70 < 82 + 88+N < 85; 80 < N <85 3

OpenStudy (anonymous):

80 < 82+88+N< 85; 70< N<85 3

OpenStudy (anonymous):

82< 80+85+N < 88; 70 < N < 85 3

OpenStudy (anonymous):

(82 + 88 + x)/3 < 85 (170 + x)/3 < 85 170 + x < 255 x < 255 - 170 x < 85 (82 + 88 + x)/3 > 80 (170 + x)/3 > 80 170 + x > 240 x > 240 - 170 x > 70 70 < x < 85 first answer choice

OpenStudy (anonymous):

Really! Thank you!!! You are saving my life!